OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. –2x – y = –7 2x – y = –7 2x – y = 11 –2x + y = –11

Parth (parthkohli):

\[y- y_1 = m(x-x_1) \]

Parth (parthkohli):

First plug-in values, then convert that equation into standard form.

OpenStudy (anonymous):

EQUATIONS TO MEMORIZE slope formula \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept \[y=mx+b\]

Parth (parthkohli):

\[ m = SLOPE = 2\]\[ (x_1,y_1) = (-1,5)\]

OpenStudy (anonymous):

so y-5=3(x-2)